Wednesday, April 25, 2018

SCS NMR Spectrometer Status Now

Updated 25 April 2018 (DLO):
Working Status of Spectrometers.  M - F, 9 a.m. to 5:30 p.m., just walk in.
After business hours, use your i-Card (swipe several times as needed).

Use Mnova Version 11.0.1 Only.

RAL Spectrometers (Check ChemFOM for Availability):
  • Carver-Bruker 500 CryoProbe (CB500) Down.  Working on problem.
    • Lift gas flow rate slightly increased.
    • Ask a staff member for account registration, then a trained person in your group for automation instruction.  Be careful with 2D NMR set-up.
  • U500 - OK, but no spinning.
  • UI400 - OK. H, B-11, F, P only.  No C-13.
  • UI500NB - OK.  Shimming to be improved soon.  Standard probe; skip tuning and matching.  OK for all 1D and 2D NMR, but calibrate pw90 and T1 for protons using posted modification (very minor).
  • VXR500 - OK, H & P only; Standard probe broken.  VT -60 to +80 C.
  • SUNDS1 and Printers - OK.
Noyes Lab Spectrometers: 
  • UI300WB- OK; Reservations via Andre Sutrisno. 
  • UI600 - OK; Reservations via Lingyang Zhu.   
Chem Life Spectrometer:
  • VNS750NB - OK; Reservations via Lingyang.  Currently in liquids configuration.